OpenStudy (anonymous):

decide whether the pair of lines is parallel, perpendicular or neither 3x + 4y = 5 4x + 3y = 9

OpenStudy (anonymous):

If the slope is the same, parallel, if it is the opposite reciprocal, pependicular

OpenStudy (anonymous):

neither

OpenStudy (anonymous):

y = -3/4x + 5/4 y = -4/3x + 3 Neither